Density of a gas is 1.27 gL^-1 at 50°c and 0.987×10^5 Pa. Claculate it's molar mass.
 

Grade:11

1 Answers

Naureen Nida
5 years ago
PV=nRT
given, d=1.27gL-1          T=50C = 323K      P=0.987×105 Pa        R=8.3145L kPa / mol.K
P(m/d)=(m/M)RT
M=dRT/P
 M= 1.27×8.315×323/0.987×105
 M=3410.69104/98700 
M= 0.03455614 g/mol